# Finite Automata Extraction: Low-data World Model Learning as Programs from Gameplay Video

### v1

@@ Results
In this paper, we proposed a framework to learn code from gameplay video. Unfortunately there are some limitations to the existing approach. In our current set up, the grid representations that were extracted for the Sprite Extraction Network occasionally returns inaccurate frames. Occasionally a sprite can be between two grid cells. In these instances, our grid representation hallucinated that there was a sprite in both of the cells. This occasionally caused the model to learn conditions that were not true. Perhaps in the future, we can tweak the Marionette Architecture to take this into account. Another aspect that the current representation does not handle the rotation of the sprites. The sprites currently are rendered in only one rotation. As noted in the case of River Raid, the grid based representation in its current state does not handle sprites with different sizes well.
@@
Retro coder also in its current state has some limitations. Since we are learning programs for individual sprites, we cannot create a GameManager program to handle sprite spawning. Additionally, the DSL does not take time into account, and can thus not make behaviours that are time dependent. For example, a certain sprite changing into a different sprite every $x$ seconds. The DSL also cannot replicate randomised behaviour.
Conclusion
@@
In this paper, we proposed a neuro symbolic approach to learn code from gameplay video. We did this by converting the frames into a symbolic grid representation. Additionally we defined a novel DSL, Retro coder for the search space of the programs. The model uses low data compared to current world models. We tested our pipeline in two domains, Pac-Man and River Raid. Our model was able to learn code for some of the sprites and showed comparable results to the existing baselines.

### v2

@@ 8. Limitations
In this paper, we proposed a framework to learn code from a gameplay video. Our results indicate that we can learn more general programs than a prior DSL-based world model, and can outperform a neural network world model on this low amount of data. However, there are still some limitations to our existing approach. In our current set up, the grid representations that were extracted from the Marionette variant occasionally return inaccurate frames. For example, occasionally a sprite can be between two grid cells. In these instances, our grid representation hallucinated that there was the same sprite in both of the cells. This caused the model to learn conditions that were not true. Perhaps in the future, we can tweak the Marionette Architecture to take this into account. Another aspect that the current representation does not handle is the rotation of the sprites. The sprites currently are rendered in only one rotation. As noted in the case of River Raid, the grid based representation in its current state also does not handle sprites with different sizes well.
@@
In its current state, Retro Coder also has some limitations. Since we learn programs for individual sprites, we cannot create a central manager or control program to handle sprite spawning. Additionally, the DSL does not take time into account, and can thus not make behaviours that are time dependent. For example, a certain sprite changing into a different sprite every $x$ seconds. The DSL also cannot replicate randomized behaviour.
9. Conclusion
@@
In this paper, we proposed a neuro-symbolic approach to learn a world model as code from gameplay video: Finite Automata Extraction (FAE). We did this by training a model to convert the video frames into a symbolic grid representation. Additionally we defined a novel DSL, Retro Coder as the search space of the world model programs. FAE uses lower data compared to current neural network world models. We tested our pipeline in two domains, Pac-Man and River Raid. Our model was able to learn accurate code for some of the sprites and showed equivalent or better results compared to the existing baselines.
Acknowledgements.
This work was funded by the Canada CIFAR AI Chairs Program. We acknowledge the support of the Alberta Machine Intelligence Institute (Amii).